19. Who goes to the convention? A club has 30 student members and 10 faculty members. The students are Abel Carson Ghosh Chen Griswold Jones Neyman Shaw DavidHein Deming Hernandez Klotz Pearl Elashoff Holland Liu Fisher Huber Miranda Reinmann Jimenez Moskowitz Santos Kim OBrien Thompson Utts Varga Potter The faculty members are Andrews Fernandez Kim BesicovitchGupta Lightman Phillips Yang Moore West The club can send 4 students and 2 faculty mem bers to a convention. It decides to choose those who will go by random selection. Describe a method for using Table D to select a stratified random sample of 4 students and 2 faculty. Then use line 123 to select the sample.
media%2Ff6d%2Ff6dcd376-f484-4225-9e22-e8
Answer #1

Here we can form two stratas namely, the student population and the faculty population. We are to choose 4 from the first strata and 2 from the other. We can proceed, by assigning numbers to each and every student. This we do as follows:

0 1 2 3 4
5 6 7 8 9
10 11 12 13 14
15 16 17 18 19
20 21 22 23 24
25 26 27 28 29

Selection Rule for students:

  • Suppose R is a two-digit random number drawn.
  • If R \leq 29 we select the R-th student.
  • If R \geq 90 then we do not select any student and draw another random number.
  • If  30\leq R \leq 89, then we find the remainder when R is divided by 30. And the remainder is used to select.

Also we assign numbers to the faculties:

0 1 2 3 4
5 6 7 8 9

Selection Rule for faculties:

  • Suppose R' is a two digit random number drawn.
  • If R' \leq 9 we select the R-th faculty.
  • If  10 \leq R' \leq 99, then we find the remainder when R' is divided by 10. And the remainder is used to select.

First of all we draw 4 random numbers from the table. And corresponding to the number drawn we select the Student given the same number as in the table.

Random Number Remainder when R is divided by 30 Student Chosen
54 24 Utts
58 28 Potter
08 - Moskowitz
15 - David

Continuing, we draw random numbers to select two faculties:

Random Number Remainder when R is divided by 10 Faculty Selected
7 - Lightman
27 7 Reject Number (Already selected)
10 0 Andrews

Note: The numbers are designated as according to the arrangement in the question.